Make New Entry,
Make Followup Entry
User name xqian
Log entry time 07:56:08 on August21,2009
Entry number 285092
keyword=Owl Shift Summary
2:30 am, MCC called, the CW beam will be delivered soon.
3:00 am asked MCC to do a HARP scan
3:15 am MCC called (finish the HARP scan on H03A and H03B)
3:20 am We found the harp scan date does not match with the current time. We called MCC for a new HARP scan on H03A and H03B.
3:25 am MCC called, the harp analysis still shows the old results. We called the RC and decide to wait till morning.
3:45 am We asked MCC to do harp scan for C07A and C07B. We do not see any peaks there. So we will wait till morning.
4:30 MCC called. They will send tune beam. We stop the current run 11503.
6:00 beam will be back shortly start a new run